Let A = [–1, 1].  
 discuss whether the following functions defined on A are one-one, onto or bijective: 
  k(x) = x2

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

k (x) = x^2
Let k (x1) = k (x2)
x1^2= x2^2
x1= ± x2
Therefore, k (x) is not one-one.

Thanks